Three cocks A, B and C bake a special kind of cake with respective probabilities as 0.02, 0.03 and 0.05 it fails to rise. In the restaurant where they work, A bakes 50 % of these cakes, B bakes 30 % of these cakes and C bakes 20 % of these cakes. What proportion of failure caused by A, B and C?



1
Expert's answer
2021-10-06T17:05:14-0400

Solution:

The proportion of failure caused by A"=50\% \times0.02=\dfrac{50}{100}\times 0.02=\dfrac1{100}"

The proportion of failure caused by B"=30\% \times0.03=\dfrac{30}{100}\times 0.03=\dfrac9{1000}"

The proportion of failure caused by C"=20\% \times0.05=\dfrac{20}{100}\times 0.05=\dfrac1{100}"